OpenStudy (anonymous):

jessica and castel each improved their yards by planting daylilies and ornamental grass. they bought their supplies from the same store. jessica spent $47 dollars on 2 daylillies and 3 bunches of ornamental grass. Castel spent $87 on 6 daylilies and 3 bunches of ornamental grass. find the cost of one daylily and the cost of one bunch of ornamental grass.

OpenStudy (anonymous):

This is two equations two variables. Let x be the cost of 1 bunch daylillies and y be the cost of 1 bunch of ornamental grass. So 2x + 3y = 47 and 6x + 3y = 87 Do you understand so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Put those on top of each other (one above the other) The idea is to get rid of one of the letters. so we should have a 3y over a 3y. If the bottom one was -3y we could add them up and they would go away. So. let's change the sign of every term on the bottom. 2x + 3y = 47 -6x - 3y = -87 OK so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So add straight down and you get -4x = -40 Divide both sides by -4 and you get x = 10 Plug that into one of the equations and solve for y 2x + 3y = 47 2(10) + 3y = 47 20 + 3y = 47 subtract 20 from both sides 3y = 27 divide both sides by 3 y = 9 So the cost of one daylilly was $10 (because x = 10) The cost of one ornamental was $9 (because y = 9)

OpenStudy (anonymous):

oo okay

OpenStudy (anonymous):

thanku